40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 26 02:15:36 2020, gautam, Update, LSC, Lock acquisition portal entry 
    Reply  Tue May 26 02:31:00 2020, gautam, Update, LSC, Lock acquisition sequence PRFPMIlock_1274418200_1274418550.pdf
       Reply  Wed May 27 20:14:51 2020, Koji, Update, LSC, Lock acquisition sequence 
    Reply  Tue May 26 02:37:19 2020, gautam, Update, LSC, DARM loop measurement and fitting DARM_TF.pdfDARM_TF_breakdown.pdf
    Reply  Tue May 26 03:01:35 2020, gautam, Update, LSC, CARM loop  CARM_OLTF.pdf
       Reply  Wed Jun 3 01:46:14 2020, gautam, Update, LSC, CARM loop  CM_loop_topology.pdfCARM_TFs.pdfCARM_OLTF.pdfCARM_xover.pdfCARM_OLG_evolution.pdf
    Reply  Tue May 26 03:06:59 2020, gautam, Update, LSC, PRFPMI sensing matrix PRFPMI_20200524sensMat.pdfPRFPMI_20200524sensMatHistograms.pdf
    Reply  Tue May 26 03:26:58 2020, gautam, Update, LSC, Preliminary noise budget PRFPMI_NB.pdf
    Reply  Tue May 26 14:32:44 2020, gautam, Update, LSC, Arm transmission RIN armRIN.pdf
       Reply  Wed May 27 19:36:33 2020, Koji, Update, LSC, Arm transmission RIN 
          Reply  Thu May 28 18:36:45 2020, gautam, Update, LSC, Arm transmission RIN PRFPMIcorner_ASC_PIT_1274419354_1274419654.pdfPRFPMIcorner_ASC_YAW_1274419354_1274419654.pdfPRFPMIcorner_ASC_coherence_1274419354_1274419654.pdf
             Reply  Fri May 29 00:34:57 2020, rana, Update, LSC, Arm transmission RIN 
             Reply  Wed Jun 3 02:14:32 2020, gautam, Update, ASC, PRC ASC improves arm transmission RIN PRC_ASCsignals.pdfarmRIN_PRC_ASC.pdfPRFPMIcorner_ASC_PIT_1275190251_1275190551.pdfPRFPMIcorner_ASC_YAW_1275190251_1275190551.pdfPRFPMIcorner_ASC_coherence_1275190251_1275190551.pdf
                Reply  Fri Jun 19 16:30:09 2020, gautam, Update, ASC, Some thoughts about ASC sensingResponse.pdfsensingResponse_torque.pdf
                   Reply  Tue Jul 7 14:06:10 2020, gautam, Update, ASC, Some more thoughts about ASC ITM_OL_DCcoupling.png
    Reply  Tue May 26 16:00:06 2020, gautam, Update, LSC, Power buildup diagnostics PRFPMIcorner_DC_1274419354_1274419654.pdfPRFPMIcorner_SB_1274419354_1274419654.pdf
       Reply  Wed May 27 17:41:57 2020, Koji, Update, LSC, Power buildup diagnostics 
          Reply  Wed Jun 3 02:08:00 2020, gautam, Update, LSC, Power buildup diagnostics PRFPMIcorner_DC_1275190251_1275190551.pdfPRFPMIcorner_SB_1275190251_1275190551.pdf
    Reply  Wed Jun 3 01:34:53 2020, gautam, Update, LSC, Lock acquisition update portal 
       Reply  Wed Jun 3 03:29:26 2020, Koji, Update, LSC, Lock acquisition update portal 
          Reply  Wed Jun 3 11:40:56 2020, gautam, Update, LSC, Lock acquisition update portal 
             Reply  Wed Jun 3 18:49:47 2020, gautam, Update, LSC, PRG and CARM signal sign armCavReflectivities.pdfIFOreflectivities.pdfPDHerrSigs.pdfPRGvsLoss_finesse.pdf
Message ID: 15350     Entry time: Tue May 26 02:37:19 2020     In reply to: 15348
Author: gautam 
Type: Update 
Category: LSC 
Subject: DARM loop measurement and fitting 

Summary:

In order to estimate the free-running DARM displacement noise, I measured the DARM OLTF using the usual IN1/IN2 prescription. The measured data was then used to fit some model paramters for a loop model that can be used over a larger frequency range.

Details:

  • Attachment #1 shows an overlay of the measured and modelled TFs.
  • Attachment #2 shows the various components that went into building up this model. 
    • The digital AA and AI filter coefficients were taken from the RTCDS code.
    • The analog AA and AI filter zpks were taken from here and here respectively.
    • CDS filters taken from the banks enabled. The 20Hz : 0Hz z:p filter in the CARM_B path is also accounted for, as have the violin-mode notches.
    • Pendulum TF is just 1/f^2, the overall scaling is unimportant because it will be fitted (in combination with the overall scaling uncertainty on the DC optical gain), but I used a value of 10 nm/f^2 which should be in the right ballbark.
    • The optical gain includes the DARM pole at ~4.5 kHz for this config.
  • With all these components, to make the measurement and fit line up, I added two free parameters - an overall gain, and a delay. 
    • NLSQ minimizer was used to find the best-fit values for these parameters.
    • I'm not sure what to make of the relatively large disagreement between measurement and model below 100 Hz - I'm pretty sure I got all the CDS filters included...
    • Moreover, I don't have a good explanation for why the best-fit delay is 400 us. One RTCDS clock cycle is onyl 60 us, and even with an extra clock cycle for the RFM transfer, I still can't get up to such a high delay...

In summary, the UGF is ~150 Hz and phase margin is ~30 deg. This loop would probably benefit from some low-pass filter being turned on.

Attachment 1: DARM_TF.pdf  126 kB  Uploaded Tue May 26 03:39:51 2020  | Hide | Hide all | Show all
DARM_TF.pdf
Attachment 2: DARM_TF_breakdown.pdf  115 kB  Uploaded Tue May 26 03:39:58 2020  | Show | Hide all | Show all
ELOG V3.1.3-